OpenStudy (anonymous):

Which of the following equations describes the line shown below? Check all that apply. Which of the following equations describes the line shown below? Check all that apply.(-3,-6) (-1,8)

OpenStudy (redohawk):

more info please :)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

A. y+6=7(x+3) B. y-6=7(x-3) C. y-8=7(x+1) D. y+8=7(x-1) E. y-8=7/2(x+1) F. y+8=7/2(x-1)

OpenStudy (anonymous):

@RedoHawk

OpenStudy (redohawk):

B, and D, and F not sure though 100%

OpenStudy (anonymous):

kk